A manufacturer has a monthly fixed cost of $42,500 and a production cost of $6 for each unit produced. The product sells for $11/unit.
(a) What is the cost function?
C(x)
=

(b) What is the revenue function?
R(x)
=

(c) What is the profit function?
P(x)
=

(d) Compute the profit (loss) corresponding to production levels of 6,000 and 11,000 units.
P(6,000)
=
P(11,000)
=

Answer by mathsolverplus(88)

C(x) = 42500+6x
R(x) = 11x
P(x) = 11x-(425000+6x) = 5x-42500

P(6000)=5(6000)-42500= -12500
P(11000)=5(11000)-42500= 12500



If this answer has helped you and you need help on additional problems, please follow us on Instagram or message us via social media! We also help with SAT Math Questions!

[Best way to contact us] Instagram (@mathsolver.plus): https://www.instagram.com/mathsolver.plus/
Twitter (@MathSolverPlus): https://twitter.com/MathSolverPlus
Snapchat (@mathsolverplus)
Email: mathsolverplus@gmail.com